We know that 10% people who purchase cars buy manual transmissions and 90% people purchase automatics.Three people purchase a car. find the probability distribution for the number who purchase automatics

1
Expert's answer
2022-03-16T08:44:10-0400

Let X be a random variable representing the number who purchase automatics, then X ~ Bin(3, 0.9)

"P(X=0)={3 \choose 0}*0.9^0*0.1^3=0.001"

"P(X=1)={3 \choose 1}*0.9^1*0.1^2=0.027"

"P(X=2)={3 \choose 2}*0.9^2*0.1^1=0.243"

"P(X=3)={3 \choose 3}*0.9^3*0.1^0=0.729"